Problem 21-03A Suttons Electronic Repair Shop has budgeted the following time and material for 2020. SUTTONS ELECTRONIC REPCompute the material loading charge percentage for Suttons Electronic Repair Shop for the year 2020. Material loading percen
Answer #1
1
Budgeted cost per labor hour 31.00 =139500/4500
Add: Profit margin 14.00
Labour rate 45.00 per hour
2
Material loading charges 58.00% =59160/102000
Add: Profit on parts 30.00%
Material loading percentage 88.00%
3
Labour charges 270 =6*45
Materials parts 400
Material loading charges 352 =400*88%
Total price of labor and material 1022